OpenStudy (anonymous):

Solve the equation. 7(2b - 5) + 3 =10 I will give a medal!

OpenStudy (opcode):

Do you know how to distribute?

OpenStudy (anonymous):

yes, but I am a beginner and need the answer quick!

OpenStudy (opcode):

@rosie07 we cannot just give out answers. 7(2b - 5) Distributing becomes: 14 b - 35

OpenStudy (opcode):

That gives us: 14 b - 35 + 3 = 10 Combine like terms: 14 b - 32 = 10

OpenStudy (opcode):

@rosie07 Do you follow?

OpenStudy (anonymous):

yep

OpenStudy (opcode):

Okay now we need to isolate b. 14b - 32 = 10 14b = 10 + 32 14b = 42 Divide both sides by 14: b = 3